# Split Gibbs Discrete Diffusion Posterior Sampling

### v2

@@ 4.2 Controlled Study using Synthetic Data
We draw 10k independent samples with each algorithm and compute the frequency map for the first two dimensions of ${\mathbf{x}}$. We measure the Hellinger distance and total variation distance between the empirical distribution and the ground truth posterior distribution. As shown in Table 2, our method achieves substantially more accurate posterior sampling compared to baseline methods. We include visualization of the generated samples in Appendix D.3 (Fig. 8) by projecting samples onto the first two dimensions and computing their heatmaps. When the number of dimensions is small, e.g., $D=2$, all methods approximate the true posterior distribution well. However, as the dimensionality increases, DPS, SMC, and SVDD-PM struggle to capture the posterior and tend to degenerate to the prior distribution. While SGDD consistently outperforms baseline methods in terms of Hellinger distance and total variation distance, it also shows a slight deviation from the posterior distribution as the dimensionality increases.
@@ 4.3 Guided DNAs Generation
We include more baseline methods that require additional training for this task, including a discrete diffusion fine-tuning method, DRAKES [45], classifier guidance (CG) for discrete diffusion models [32], and twisted diffusion sampler (TDS) [47] which uses CG as the proposal function for sequential Monte Carlo sampling.
@@
[…]
SVDD-PM $5.41$ $5.08_{\pm 1.51}$ $49.9$$\textbf{241}_{\pm 47}$
SMC $4.15$ $3.99_{\pm 1.49}$ $39.9$$259_{\pm 13}$
CG $2.90$ $2.76_{\pm 0.94}$ $\ \ 0.0$$265_{\pm 5\hskip 5.0pt}$
TDS $4.64$ $4.62_{\pm 1.09}$ $45.3$$257_{\pm 9\hskip 5.0pt}$
DRAKES w/o KL $6.44$ $6.28_{\pm 0.76}$ $82.5$$281_{\pm 6\hskip 5.0pt}$
DRAKES w/ KL $5.61$ $5.24_{\pm 1.39}$ $92.5$$264_{\pm 9\hskip 5.0pt}$
SGDD $(\beta=30)$ $8.82$ $\underline{8.69}_{\pm 1.12}$ $91.7$$\underline{246}_{\pm 41}$
@@
Results. We generate $640$ DNA samples using different algorithms and compute the median and average Pred-Activity of the generated sequences. As shown in Table 3, our method achieves the highest Pred-Activity (reward values) compared to existing baselines, including 42% higher median reward compared to the previous state-of-the-art method.
Moreover, as SGDD samples from the posterior distribution, $p({\mathbf{x}})\exp(\beta r({\mathbf{x}}))/Z$, the guidance strength $\beta$ controls the strength of guidance imposed in the sampling algorithm. We instantiate SGDD with $\beta$ ranging from $3$ to $100$ and calculate the Pred-Activity and Log-likelihood of the generated samples. We plot the distribution of the samples’ Pred-Activity with various $\beta$ in Fig. 3. In general, using a larger $\beta$ results in higher Pred-Activity (reward), at the expense of a lower likelihood as the target distribution deviates further from the prior distribution.
@@ 4.4 Discrete Image Inverse Problems
[…]
@@
Results. As shown in Table 4, SGDD outperforms baseline methods by a large margin in both XOR and AND tasks. We present the samples generated by SGDD for the XOR task in Section D.2. The reconstructed samples are visually consistent with the underlying ground truth signal, which explains the high class accuracy of our method shown in Table 4.
Sample diversity. Furthermore, we demonstrate that SGDD generates diversified samples when the measurement ${\mathbf{y}}$ is sparse. For example, when a digit is masked with a large box, as shown in Fig. 3, the measurement lacks sufficient information to fully recover the original digit. In this scenario, SGDD generate samples from multiple plausible modes, including digits 1,4,7 and 9. This highlights the ability of our method to produce diverse samples from the posterior distribution while preserving consistency with the measurement.
Higher dimensional images. To demonstrate the scalability of SGDD, we further evaluate SGDD on a higher-dimensional dataset, FFHQ 256 [21] in Section D.1, where we solve the super-resolution task by operating in the discrete latent space of a pretrained vector-quantized diffusion model.
@@ 4.5 Monophonic Music
We conduct experiments on monophonic songs from the Lakh pianoroll dataset [11]. Following the preprocessing steps in [4], we obtain monophonic musical sequences of length $D=256$ and vocabulary size $N=129$. The ordering of musical notes is scrambled to destroy any ordinal structure. Therefore, the likelihood function for this task is $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}})\|_{0}/\sigma_{{\mathbf{y}}})$. This $\ell_{0}$ norm makes this inverse problem ill-posed, in stark contrast to most reward-guided generation tasks, where the reward functions are much smoother. We evaluate our method on an inpainting task, where the forward model randomly masks $\gamma=40\%,60\%$ of the notes. As in Campbell et al. [4], we use the Hellinger distance of histograms and the proportion of outlier notes as metrics to evaluate our method. We run experiments on 100 samples in the test set and report the quantitative results in Table 5.  SGDD successfully completes monophonic music sequences in a style more consistent with the given conditions, achieving a 2x reduction in Hellinger distance compared to SVDD-PM.
@@ Appendix A Theory
For two probability mass function $\mu$ and $\pi$, we discuss their KL divergence,
$\mathsf{KL}(\mu\|\pi):=\mathbb{E}_{{\mathbf{x}}\sim\mu}\left[\ln\frac{\mu}{\pi}({\mathbf{x}})\right].$ (16)
@@
In continuous state spaces, the relative Fisher information criterion has been used to derive general first-order guarantees for non-log-concave sampling [2]. This line of analysis has been adapted to provide theoretical insights for posterior sampling methods using diffusion models [41, 48]. The formula in Eq. 17 represents a discrete state space analogue of Fisher information. We refer interested readers to [3, 10, 17] for more discussions on this topic and the relation between KL and Fisher information in the discrete state space. Our analysis in this paper adapts these techniques and provides general first-order guarantees for posterior sampling with discrete diffusion models.
@@ A.1 Lemmas
$\partial_{t}\mathsf{KL}(\mu_{t}\|\pi_{t})\leq-\frac{1}{2}\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})+\frac{2}{c}\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\tilde{\bm{Q}}_{t}\mu_{t}),$ (19)
where $\mathcal{L}_{{\bm{Q}}}(\mu_{t},\partial_{t}\mu_{t})\geq 0$ is the Lagrangian defined as
$\mathcal{L}_{{\bm{Q}}}(\mu_{t},\partial_{t}\mu_{t})=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\left[\langle\varphi,\partial_{t}\mu_{t}\rangle-\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}\in\mathcal{X}}\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}^{[j,i]}\exp(\varphi({\mathbf{x}}_{j})-\varphi({\mathbf{x}}_{i}))\right].$ (20)
@@
$\partial_{t}\mathsf{KL}(\mu_{t}\|\pi_{t})$ $=\partial_{t}\sum_{{\mathbf{x}}_{i}}\mu_{t}({\mathbf{x}}_{i})\log\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}=\sum_{{\mathbf{x}}_{i}}\left(\partial_{t}\mu_{t}({\mathbf{x}}_{i})\log\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}-\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}\partial_{t}\pi_{t}({\mathbf{x}}_{i})\right)$
$=\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}\left(\tilde{{\bm{Q}}}_{t}^{[i,j]}\mu_{t}({\mathbf{x}}_{j})\log\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}-\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}{\bm{Q}}_{t}^{[i,j]}\pi_{t}({\mathbf{x}}_{j})\right)$
$=\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}{\bm{Q}}_{t}^{[i,j]}\pi_{t}({\mathbf{x}}_{j})\left(\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}\log\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}-\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}\right)-\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}({\bm{Q}}_{t}^{[i,j]}-\tilde{{\bm{Q}}}_{t}^{[i,j]})\mu_{t}({\mathbf{x}}_{j})\log\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}$
Using the fact that $\sum_{{\mathbf{x}}_{i}}{\bm{Q}}_{t}^{[i,j]}=0$ and the definition in Eq. 17, we have the equality
$\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}{\bm{Q}}_{t}^{[i,j]}\pi_{t}({\mathbf{x}}_{j})\left(\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}\log\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}-\frac{\mu_{t}({\mathbf{x}}_{i})}{\pi_{t}({\mathbf{x}}_{i})}\right)=-\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})$
Using the relation $\partial_{t}\mu_{t}=\tilde{\bm{Q}}_{t}\mu_{t}$ then leads to
$\partial_{t}\mathsf{KL}(\mu_{t}\|\pi_{t})=-\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})-\underbrace{(\log\frac{\mu_{t}}{\pi_{t}})^{T}({\bm{Q}}_{t}-\partial_{t})\mu_{t}}_{\textit{error term}}.$ (21)
@@
$\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\partial_{t}\mu_{t})$ $=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}[\langle\varphi,\partial_{t}\mu_{t}\rangle-(e^{-\varphi}\mu_{t})^{T}{\bm{Q}}_{t}^{T}e^{\varphi}]$
$=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}[\langle\varphi,\partial_{t}\mu_{t}-{\bm{Q}}_{t}\mu_{t}\rangle-\underbrace{\mu_{t}^{T}(e^{-\varphi}{\bm{Q}}_{t}^{T}e^{\varphi}-{\bm{Q}}_{t}^{T}\varphi)}_{\textit{denotedas}\ \tilde{\mathcal{H}}(\mu_{t},\varphi)}]$ (22)
@@
$\langle\varphi,\partial_{t}\mu_{t}-{\bm{Q}}_{t}\mu_{t}\rangle\leq\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\partial_{t}\mu_{t})+\tilde{\mathcal{H}}(\mu_{t},\varphi).$ (23)
If we choose $\varphi=\log\frac{\mu_{t}}{\pi_{t}}$, the inequality above gives a bound to the error term in Eq. 21. Moreover, it is easy to verify that
~$\tilde{\mathcal{H}}(\mu_{t},\log\frac{\mu_{t}}{\pi_{t}})=\mathsf{FI}(\mu_{t}\|\pi_{t}).$ (24)
In fact, when the space $\mathcal{X}$ is continuous, choosing $\varphi({\mathbf{x}})=\log\frac{\mu_{t}}{\pi_{t}}({\mathbf{x}})$ exactly recovers Lemma A.4 in [48]. However, as pointed out in [16], it is necessary to consider a rescaled $\varphi=\lambda\log\frac{\mu_{t}}{\pi_{t}}$ with $\lambda\in(0,1)$ to derive a bound in finite space $\mathcal{X}$.
@@
~$\tilde{\mathcal{H}}(\mu_{t},\lambda\varphi)\leq\frac{\lambda^{2}}{c}\tilde{\mathcal{H}}(\mu_{t},\varphi).$ (25)
Plugging in $\varphi=\lambda\log\frac{\mu_{t}}{\pi_{t}}$ in Eq. 23, we have
$\langle\lambda\log\frac{\mu_{t}}{\pi_{t}},\partial_{t}\mu_{t}-{\bm{Q}}_{t}\mu_{t}\rangle$ $\leq\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\partial_{t}\mu_{t})+\tilde{\mathcal{H}}(\mu_{t},\lambda\log\frac{\mu_{t}}{\pi_{t}})$ (26)
$\leq\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\partial_{t}\mu_{t})+\frac{\lambda^{2}}{c}\tilde{\mathcal{H}}(\mu_{t},\log\frac{\mu_{t}}{\pi_{t}})$ (27)
Combine this with Eq. 21,
$\partial_{t}\mathsf{KL}(\mu_{t}\|\pi_{t})$ $\leq-\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})+\frac{1}{\lambda}\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\partial_{t}\mu_{t})+\frac{\lambda}{c}\tilde{\mathcal{H}}(\mu_{t},\log\frac{\mu_{t}}{\pi_{t}})$
$=-(1-\frac{\lambda}{c})\mathsf{FI}_{{\bm{Q}}_{t}}(\mu_{t}\|\pi_{t})+\frac{1}{\lambda}\mathcal{L}_{{\bm{Q}}_{t}}(\mu_{t},\partial_{t}\mu_{t}).$ (28)
@@
When the learned score function $s_{\theta}$ satisfies Assumption (i),
@@
$=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}\in\mathcal{X}}\left[\mu_{t}({\mathbf{x}}_{i})\tilde{\bm{Q}}_{t}^{[j,i]}(1+\varphi({\mathbf{x}}_{j})-\varphi({\mathbf{x}}_{i}))-\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}e^{\varphi({\mathbf{x}}_{j})-\varphi({\mathbf{x}}_{i})}\right]$ ($\mathbf{1}^{T}\tilde{\bm{Q}}_{t}{\mathbf{u}}=0$ for any ${\mathbf{u}}$)
$=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}\Big{(}\frac{\tilde{\bm{Q}}_{t}^{[j,i]}}{{\bm{Q}}_{t}^{[j,i]}}(1-z_{ij})-e^{-z_{ij}}\Big{)}\right]+\sum_{{\mathbf{x}}_{i}}\mu_{t}({\mathbf{x}}_{i})(\tilde{{\bm{Q}}}_{t}^{[i,i]}-{\bm{Q}}_{t}^{[i,i]})$ ($z_{ij}=\varphi({\mathbf{x}}_{i})-\varphi({\mathbf{x}}_{j})$)
$\leq\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}\cdot\sup_{z_{ij}\in\mathbb{R}}\Big{(}\frac{\tilde{\bm{Q}}_{t}^{[j,i]}}{{\bm{Q}}_{t}^{[j,i]}}(1-z_{ij})-e^{-z_{ij}}\Big{)}\right]+\epsilon\sum_{{\mathbf{x}}_{i}}\mu_{t}({\mathbf{x}}_{i})|{\bm{Q}}_{t}^{[i,i]}|.$ (30)

### v3

@@ 4.2 Controlled Study using Synthetic Data
We draw 10k independent samples with each algorithm and compute the frequency map for the first two dimensions of $\mathbf{x}$}. We measure the Hellinger distance and total variation distance between the empirical distribution and the ground truth posterior distribution. As shown in Table˜2, our method achieves substantially more accurate posterior sampling compared to baseline methods. We include visualization of the generated samples in Appendix D.3 (Fig.˜8) by projecting samples onto the first two dimensions and computing their heatmaps. When the number of dimensions is small, e.g., $D=2$, all methods approximate the true posterior distribution well. However, as the dimensionality increases, DPS, SMC, and SVDD-PM struggle to capture the posterior and tend to degenerate to the prior distribution. While SGDD consistently outperforms baseline methods in terms of Hellinger distance and total variation distance, it also shows a slight deviation from the posterior distribution as the dimensionality increases.
@@ 4.3 Guided DNAs Generation
We include more baseline methods that require additional training for this task, including a discrete diffusion fine-tuning method, DRAKES [48], classifier guidance (CG) for discrete diffusion models [32], and twisted diffusion sampler (TDS) [50] which uses CG as the proposal function for sequential Monte Carlo sampling. We further include a Feynman-Kac diffusion sampler [33] with MAX potential function as well.
@@
Pred-Activity (median) $\uparrow$ Pred-Activity (average) $\uparrow$ ATAC    Acc $\%$ $\uparrow$ Log-likelihood (average) $\uparrow$
SVDD-PM $5.41$ $5.08_{\pm 1.51}$ $49.9$ -$\textbf{241}_{\pm 47}$
SMC $4.15$ $_{\pm 1.49} 39.939.9$ -$259_{\pm 13}$
CG $2.90$ $2.76_{\pm 0.94}$ 0.0\ \ 0.0 -$265_{\pm 5\hskip 5.0pt}$
TDS $4.64$ $4.62_{\pm 1.09}$ $45.3$ -$257_{\pm 9\hskip 5.0pt}$
FK-MAX $4.90$ $4.79_{\pm 1.02}$ $47.1$ $243_{\pm 9\hskip 5.0pt}$
DRAKES w/o KL $6.44$ $6.28_{\pm 0.76}$ $82.5$ -$281_{\pm 6\hskip 5.0pt}$
DRAKES w/ KL $5.61$ $5.24_{\pm 1.39}$ $92.5$ -$264_{\pm 9\hskip 5.0pt}$
SGDD $(\beta=30)$ $8.82$ $\underline{8.69}_{\pm 1.12}$ $91.7$ -$\underline{246}_{\pm 41}$
@@
Results. We generate $640$ DNA samples using different algorithms and compute the median and average Pred-Activity of the generated sequences. As shown in Table˜3, our method achieves the highest Pred-Activity (reward values) compared to existing baselines, including 42% higher median reward compared to the previous state-of-the-art method.
Moreover, as SGDD samples from the posterior distribution, $p({\mathbf{x}})\exp(\beta r({\mathbf{x}}))/Z$, the guidance strength $\beta$ controls the strength of guidance imposed in the sampling algorithm. We instantiate SGDD with $\beta$ ranging from $3$ to $100$ and calculate the Pred-Activity and Log-likelihood of the generated samples. We plot the distribution of the samples’ Pred-Activity with various $\beta$ in Fig.˜3. In general, using a larger $\beta$ results in higher Pred-Activity (reward), at the expense of a lower likelihood as the target distribution deviates further from the prior distribution.
@@ 4.4 Discrete Image Inverse Problems
PSNR $\uparrow$ Accuracy ($\%$) $\uparrow$ PSNR $\uparrow$ Accuracy ($\%$) $\uparrow$
@@
Results. As shown in Table˜4, SGDD outperforms baseline methods by a large margin in both XOR and AND tasks. We present the samples generated by SGDD for the XOR task in Section˜D.2. The reconstructed samples are visually consistent with the underlying ground truth signal, which explains the high class accuracy of our method shown in Table˜4.
Sample diversity. Furthermore, we demonstrate that SGDD generates diversified samples when the measurement $\mathbf{y}$} is sparse. For example, when a digit is masked with a large box, as shown in Fig.˜3, the measurement lacks sufficient information to fully recover the original digit. In this scenario, SGDD generate samples from multiple plausible modes, including digits 1,4,7 and 9. This highlights the ability of our method to produce diverse samples from the posterior distribution while preserving consistency with the measurement.
Higher dimensional images. To demonstrate the scalability of SGDD, we further evaluate SGDD on a higher-dimensional dataset, FFHQ 256 [21] in Section˜D.1, where we solve the super-resolution task by operating in the discrete latent space of a pretrained vector-quantized diffusion model.
@@ 4.5 Monophonic Music
We conduct experiments on monophonic songs from the Lakh pianoroll dataset [11]. Following the preprocessing steps in [4], we obtain monophonic musical sequences of length $D=256$ and vocabulary size $N=129$. The ordering of musical notes is scrambled to destroy any ordinal structure. Therefore, the likelihood function for this task is $p({\mathbf{y}}|{\mathbf{x}})\ \propto\ \exp(-\|{\bm{G}}({\mathbf{x}})-{\mathbf{y}})\|_{0}/\sigma_{{\mathbf{y}}})$. This $\ell_{0}$ norm makes this inverse problem ill-posed, in stark contrast to most reward-guided generation tasks, where the reward functions are much smoother. We evaluate our method on an inpainting task, where the forward model ${\bm{G}}(\cdot)$ randomly masks $\gamma=40\%,60\%$ of the notes. As in Campbell et al. [4], we use the Hellinger distance of histograms and the proportion of outlier notes as metrics to evaluate our method. We run experiments on 100 samples in the test set and report the quantitative results in Table˜5.  SGDD successfully completes monophonic music sequences in a style more consistent with the given conditions, achieving a 2x reduction in Hellinger distance compared to SVDD-PM.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This research was supported in part by gifts from OpenAI, Two Sigma, and Cisco. Z.W. is supported by the Amazon AI4Science fellowship. W.C. is supported by the Kortschak Scholars Fellowship.
@@ Appendix A Theory
For two probability mass functions $\mu$ and $\pi$, we discuss their KL divergence,
$\mathsf{KL}(\mu\|\pi):=\mathbb{E}_{{\mathbf{x}}\sim\mu}\left[\log\frac{\mu}{\pi}({\mathbf{x}})\right].$ (16)
@@
In continuous state spaces, the relative Fisher information criterion has been used to derive general first-order guarantees for non-log-concave sampling [2]. This line of analysis has been adapted to provide theoretical insights for posterior sampling methods using diffusion models [43, 51]. The formula in Eq.˜17 represents a discrete state space analogue of Fisher information. We refer interested readers to [3, 10, 17] for more discussions on this topic and the relation between KL and Fisher information in the discrete state space. Our analysis in this paper adapts these techniques and provides general first-order guarantees for posterior sampling with discrete diffusion models.
@@ A.1 Lemmas
$\partial_{t}\mathsf{KL}(\pi_{t}\|\mu_{t})\leq-\frac{1}{2}\mathsf{FI}_{\tilde{\bm{Q}}_{t}}(\pi_{t}\|\mu_{t})+\frac{2}{c}\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},{\bm{Q}}_{t}\pi_{t}),$ (19)
where $\mathcal{L}_{\tilde{\bm{Q}}}(\pi_{t},\partial_{t}\pi_{t})\geq 0$ is the Lagrangian defined as
$\mathcal{L}_{\tilde{\bm{Q}}}(\pi_{t},\partial_{t}\pi_{t})=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\left[\langle\varphi,\partial_{t}\pi_{t}\rangle-\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}\in\mathcal{X}}\pi_{t}({\mathbf{x}}_{i})\tilde{\bm{Q}}^{[j,i]}\exp(\varphi({\mathbf{x}}_{j})-\varphi({\mathbf{x}}_{i}))\right].$ (20)
@@
$\partial_{t}\mathsf{KL}(\pi_{t}\|\mu_{t})$ $=\partial_{t}\sum_{{\mathbf{x}}_{i}}\pi_{t}({\mathbf{x}}_{i})\log\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}=\sum_{{\mathbf{x}}_{i}}\left(\partial_{t}\pi_{t}({\mathbf{x}}_{i})\log\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}-\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}\partial_{t}\mu_{t}({\mathbf{x}}_{i})\right)$
$=\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}\left({{\bm{Q}}}_{t}^{[i,j]}\pi_{t}({\mathbf{x}}_{j})\log\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}-\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}\tilde{\bm{Q}}_{t}^{[i,j]}\mu_{t}({\mathbf{x}}_{j})\right)$
$=\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}\tilde{\bm{Q}}_{t}^{[i,j]}\mu_{t}({\mathbf{x}}_{j})\left(\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}\log\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}-\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}\right)-\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}(\tilde{\bm{Q}}_{t}^{[i,j]}-{\bm{Q}}_{t}^{[i,j]})\pi_{t}({\mathbf{x}}_{j})\log\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}$
Using the fact that $\sum_{{\mathbf{x}}_{i}}\tilde{\bm{Q}}_{t}^{[i,j]}=0$ and the definition in Eq.˜17, we have the equality
$\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}}\tilde{\bm{Q}}_{t}^{[i,j]}\mu_{t}({\mathbf{x}}_{j})\left(\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}\log\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}-\frac{\pi_{t}({\mathbf{x}}_{i})}{\mu_{t}({\mathbf{x}}_{i})}\right)=-\mathsf{FI}_{\tilde{\bm{Q}}_{t}}(\pi_{t}\|\mu_{t})$
Using the relation $\partial_{t}\pi_{t}={\bm{Q}}_{t}\pi_{t}$ this leads to
$\partial_{t}\mathsf{KL}(\pi_{t}\|\mu_{t})=-\mathsf{FI}_{\tilde{\bm{Q}}_{t}}(\pi_{t}\|\mu_{t})-\underbrace{(\log\frac{\pi_{t}}{\mu_{t}})^{T}(\tilde{\bm{Q}}_{t}-\partial_{t})\pi_{t}}_{\textit{error term}}.$ (21)
@@
$\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},\partial_{t}\pi_{t})$ $=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}[\langle\varphi,\partial_{t}\pi_{t}\rangle-(e^{-\varphi}\pi_{t})^{T}\tilde{\bm{Q}}_{t}^{T}e^{\varphi}]$
$=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}[\langle\varphi,\partial_{t}\pi_{t}-\tilde{\bm{Q}}_{t}\pi_{t}\rangle-\underbrace{\pi_{t}^{T}(e^{-\varphi}\tilde{\bm{Q}}_{t}^{T}e^{\varphi}-\tilde{\bm{Q}}_{t}^{T}\varphi)}_{\textit{denoted as}\ \tilde{\mathcal{H}}(\pi_{t},\varphi)}]$ (22)
@@
$\langle\varphi,\partial_{t}\pi_{t}-\tilde{\bm{Q}}_{t}\pi_{t}\rangle\leq\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},\partial_{t}\pi_{t})+\tilde{\mathcal{H}}(\pi_{t},\varphi).$ (23)
If we choose $\varphi=\log\frac{\pi_{t}}{\mu_{t}}$, the inequality above gives a bound to the error term in Eq.˜21. Moreover, it is easy to verify that
$\tilde{\mathcal{H}}(\pi_{t},\log\frac{\pi_{t}}{\mu_{t}})=\mathsf{FI}_{\tilde{\bm{Q}}_{t}}(\pi_{t}\|\mu_{t}).$ (24)
In fact, when the space $\mathcal{X}$ is continuous, choosing $\varphi({\mathbf{x}})=\log\frac{\pi_{t}}{\mu_{t}}({\mathbf{x}})$ exactly recovers Lemma A.4 in [51]. However, as pointed out in [16], it is necessary to consider a rescaled $\varphi=\lambda\log\frac{\pi_{t}}{\mu_{t}}$ with $\lambda\in(0,1)$ to derive a bound in finite space $\mathcal{X}$.
@@
$\tilde{\mathcal{H}}(\pi_{t},\lambda\varphi)\leq\frac{\lambda^{2}}{c}\tilde{\mathcal{H}}(\pi_{t},\varphi).$ (25)
Plugging in $\varphi=\lambda\log\frac{\pi_{t}}{\mu_{t}}$ in Eq.˜23, we have
$\langle\lambda\log\frac{\pi_{t}}{\mu_{t}},\partial_{t}\pi_{t}-\tilde{\bm{Q}}_{t}\pi_{t}\rangle$ $\leq\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},\partial_{t}\pi_{t})+\tilde{\mathcal{H}}(\pi_{t},\lambda\log\frac{\pi_{t}}{\mu_{t}})$ (26)
$\leq\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},\partial_{t}\pi_{t})+\frac{\lambda^{2}}{c}\tilde{\mathcal{H}}(\pi_{t},\log\frac{\pi_{t}}{\mu_{t}})$ (27)
Combine this with Eq.˜21,
$\partial_{t}\mathsf{KL}(\pi_{t}\|\mu_{t})$ $\leq-\mathsf{FI}_{\tilde{\bm{Q}}_{t}}(\pi_{t}\|\mu_{t})+\frac{1}{\lambda}\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},\partial_{t}\pi_{t})+\frac{\lambda}{c}\tilde{\mathcal{H}}(\pi_{t},\log\frac{\pi_{t}}{\mu_{t}})$
$=-(1-\frac{\lambda}{c})\mathsf{FI}_{\tilde{\bm{Q}}_{t}}(\pi_{t}\|\mu_{t})+\frac{1}{\lambda}\mathcal{L}_{\tilde{\bm{Q}}_{t}}(\pi_{t},\partial_{t}\pi_{t}).$ (28)
@@
When the learned score function $s_{\theta}$ satisfies Assumption (i) and (v),
@@
$=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\sum_{{\mathbf{x}}_{i},{\mathbf{x}}_{j}\in\mathcal{X}}\left[\mu_{t}({\mathbf{x}}_{i})\tilde{\bm{Q}}_{t}^{[j,i]}(\varphi({\mathbf{x}}_{j})-\varphi({\mathbf{x}}_{i}))-\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}e^{\varphi({\mathbf{x}}_{j})-\varphi({\mathbf{x}}_{i})}\right]$ ($\mathbf{1}^{T}\tilde{\bm{Q}}_{t}{\mathbf{u}}=0$ for any $\mathbf{u}$})
$=\sup_{\varphi\in\mathcal{C}_{b}(\mathcal{X})}\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}\Big(-\frac{\tilde{\bm{Q}}_{t}^{[j,i]}}{{\bm{Q}}_{t}^{[j,i]}}z_{ij}-e^{-z_{ij}}\Big)\right]+\sum_{{\mathbf{x}}_{i}}\mu_{t}({\mathbf{x}}_{i})(\tilde{{\bm{Q}}}_{t}^{[i,i]}-{\bm{Q}}_{t}^{[i,i]})$ ($z_{ij}=\varphi({\mathbf{x}}_{i})-\varphi({\mathbf{x}}_{j})$)
$\leq\sum_{{\mathbf{x}}_{i}\neq{\mathbf{x}}_{j}}\left[\mu_{t}({\mathbf{x}}_{i}){\bm{Q}}_{t}^{[j,i]}\cdot\sup_{z_{ij}\in\mathbb{R}}\Big(-\frac{\tilde{\bm{Q}}_{t}^{[j,i]}}{{\bm{Q}}_{t}^{[j,i]}}z_{ij}-e^{-z_{ij}}\Big)\right]+\epsilon\sum_{{\mathbf{x}}_{i}}\mu_{t}({\mathbf{x}}_{i})|{\bm{Q}}_{t}^{[i,i]}|.$ (30)